# Masks and Boxes: Combining the Best of Both Worlds for Multi-Object Tracking

### v2

@@ Abstract
Multi-object tracking (MOT) involves identifying and consistently tracking objects across video sequences. Traditional tracking-by-detection methods, while effective, often require extensive tuning and lack generalizability. On the other hand, segmentation mask-based methods are more generic but struggle with tracking management, making them unsuitable for MOT. We propose a novel approach, McByte, which incorporates a temporally propagated segmentation mask as a strong association cue within a tracking-by-detection framework. By combining bounding box and mask information, McByte enhances robustness and generalizability without per-sequence tuning. Evaluated on four benchmark datasets - DanceTrack, MOT17, SoccerNet-tracking 2022, and KITTI-tracking - McByte demonstrates performance gain in all cases examined. At the same time, it outperforms existing mask-based methods. Implementation code will be provided upon acceptance.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 1: Visual output comparison between the baseline and McByte. With the mask guidance, McByte can handle longer occlusion in the crowd - see the subject with ID 54 on the output of McByte (at frame 319, on the right). Input image data from  [28]. Best seen in color.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Visual output comparison between the baseline and McByte. With the mask guidance, McByte can handle the association of an ambiguous set of bounding boxes and reduce the identity switches - see the subjects with IDs 59 and 63 on the output of McByte (at frame 459, in the middle). Input image data from [28]. Best seen in color.
@@ 1 Introduction
In this paper, we explore using a temporally propagated segmentation mask as an association cue to assess its effectiveness in MOT. We propose a novel tracking-by-detection method that combines segmentation masks and bounding boxes to improve the association between tracklets and detections. The mask is managed according to the tracklet lifespan, with usage constraints to enhance tracking performance. Since the mask propagation model is trained on a large dataset, it makes the entire tracking process more generic. Unlike existing tracking-by-detection methods, our approach does not require tuning hyperparameters for each video sequence.
We evaluate our incorporation of the mask as an association cue against a baseline tracker, showing clear benefits for MOT in our controlled approach. Our tracker is tested on four MOT datasets, achieving the highest performance from tracking-by-detection algorithms on DanceTrack [33], SoccerNet-tracking 2022 [9], and KITTI-tracking [16] (Pedestrian). On MOT17 [28], we see a small improvement over the baseline. These results highlight the advantages of using segmentation masks, eliminating the need for per-sequence hyper-parameter tuning. Visual examples of our designed tracker utilizing the mask as an association cue improving over the baseline algorithm are presented in Figs. 1 and 2.
@@
    We propose a novel approach adapting a temporally propagated object segmentation mask as a strong and powerful cue for MOT, as described in Sec. 3.3.
@@
    We design an MOT tracking algorithm substantiating the idea and incorporating the mask as an association cue between tracklets and detections, detailed in Secs. 3.1 and 3.2. The tracker overcomes the limitations of mask-based approaches by performing proper tracklet management and including other important association cues as well as the limitations of the baseline tracking-by-detection approaches, by making the tracking process more robust and generic.
@@
    We evaluate and discuss the use of the mask as an association cue with different conditions within a tracking-by-detection algorithm in the ablation study (Sec. 4.3). Further, we evaluate our tracker incorporating the mask cue on four different datasets (in Sec. 4.4), comparing it to state-of-the-art tracking-by-detection approaches and demonstrating performance gain thanks to the attentive mask usage, while not tuning per-sequence parameters.
@@ 2 Related Work
Different types of algorithms tracking multiple objects at the same time have been developed in the community. We provide a brief overview of the existing methods related to our work.
2.1 Transformer-based methods
Transformer-based methods [14, 39, 13, 37] use attention mechanisms to learn tracking trajectories and object associations from training data, following an end-to-end approach. These methods perform well on video sequences where subjects remain mostly in the scene, such as DanceTrack [33], but struggle with sequences where subjects frequently enter and exit, like in the MOT17 dataset [28], despite being designed for MOT. For this reason, we do not directly compare our method with transformer-based approaches in the main paper, though their results are listed together with the other approaches and ours in the supplementary material, Appendix B.
@@ 2.2 Tracking-by-detection methods
ByteTrack [38] is a powerful tracking-by-detection algorithm that uses the YOLOX [15] detector, associating tracklets based on the intersection over union (IoU) between bounding boxes of tracklets and new detections. It provides strong tracklet management and serves as a solid baseline for MOT. Several works have built on ByteTrack. OC-SORT [4] enhances state estimation by computing virtual trajectories during occlusion. StrongSORT [12] adds re-identification (re-ID) features, camera motion compensation, and NSA Kalman filter [11]. C-BIoU [35] extends the association process by buffering (enlarging) bounding boxes, while HybridSORT [36] adds cues like confidence modeling and height-modulated IoU alongside existing strong cues.
Although these algorithms perform well on popular MOT datasets, they are highly sensitive to parameters. ByteTrack, for instance, explicitly tunes parameters like high-confidence detection and tracklet initialization thresholds per test sequence, as noted on its GitHub page [^1] https://github.com/ifzhang/ByteTrack, see: Test on MOT17. and in the code [^2] https://github.com/ifzhang/ByteTrack/blob/main/yolox/evaluators/mot_evaluator.py, lines 146-157 (at the moment of submitting this paper). As we show in the supplementary material, in Appendix D, varying these parameters significantly impacts tracking performance. Extensions of ByteTrack [12, 4, 27, 36], sharing the same code base also rely on per-sequence parameter tuning. This tuning process is costly and impractical for larger datasets like DanceTrack [33], SoccerNet-tracking [9], and KITTI-tracking [16], as well as for generalizing across datasets. In contrast, our method, incorporating the segmentation mask as an association cue, avoids per-sequence tuning, making it more robust and generic, as we further demonstrate in Sec. 4.4.
Figure 3: Our tracking pipeline with the mask cue guidance. Mask signal is incorporated as an association cue in the tracklet-detection association steps.
@@ 2.3 Segmentation mask models
Mask temporal propagation. XMem [8] is a mask temporal propagation model for video object segmentation, based on the Atkinson-Shiffrin Memory Model [1], enabling long-term tracking of segmentation masks. Its successor, Cutie [6], improves segmentation by incorporating object encoding from mask memory and better distinguishing the object from the background. While useful for re-identifying subjects across frames, both XMem and Cutie are not suitable for MOT as they do not involve bounding boxes and can provide inaccurate mask predictions as listed in their performance on video object segmentation [8, 6]. Therefore, we propose a novel MOT algorithm that combines temporally propagated masks with bounding boxes, improving tracking performance. Our ablation studies in Sec. 4.3 show that using the mask in a controlled manner, alongside other MOT mechanisms, is more effective than relying solely on the mask signal.
Mask-based tracking systems. Segmentation mask models have already been used to build tracking systems. DEVA [7], which enhances XMem [8], proposes decoupled video segmentation and bi-directional propagation, and builds a tracking system with boxes and masks. However, MOT requires a robust tracklet management and occlusion handling system, which is not available within DEVA.
@@
Existing mask-based tracking algorithms, though trained on large datasets for extracting semantic features, fail to handle challenges like occlusion and track initiation. As shown in Sec. 4.5, their performance on MOT tasks is significantly lower than other methods. Our proposed tracker incorporates the mask as an association cue while leveraging powerful mechanisms like tracklet management and additional cues (e.g., bounding box position, motion). As we demonstrate in Sec. 4.5, our approach performs considerably better than existing mask-based systems, even when they are adapted to MOT format and evaluation.
@@ 3.1 Preliminaries
In our work, we study a temporally propagated segmentation mask as a powerful association cue for MOT. We extend the baseline algorithm [38] and combine the mask information with bounding box information to create our novel masked-cued algorithm, which we call McByte. Fig. 3 shows the overview ouf our tracking pipeline pointing where the mask is involved as an association cue.
In the next sections, we describe the creation and handling of the mask within the tracking algorithm (Sec. 3.2) and our regulated use of the mask as an association cue (Sec. 3.3).
@@ 3.2 Mask creation and handling
Since the mask as a cue has to facilitate associating a tracklet with the correct detection, it must be well correlated with the tracklets and handled accordingly. This process is originally not straightforward and thus we design the following approach of mask handling to synchronise it with the processed tracklets.
During tracking in McByte, each tracklet gets its own mask, which is propagated across frames to update the mask predictions. We use an image segmentation model to create a mask for each new tracklet, which is then passed to a temporal propagator that generates mask predictions for all tracklets during each frame. These predictions are analyzed and used in the tracklet-detection association process, as detailed in Sec. 3.3. We manage the masks in sync with the tracklet management system, creating new masks for new tracklets and removing them when a tracklet is terminated. Visual examples of the masks being used within our tracker are shown in Figs. 1 and 2.
@@ 3.3 Regulated use of the mask
A temporally propagated segmentation mask can be a strong association cue if used correctly. However, mask predictions from the temporal propagator can sometimes be incorrect, as seen in related works [8, 6] and thus unreliable, making it essential to regulate the mask’s use as a cue.
In the association process, we update the cost matrix entries using the mask, particularly in cases of ambiguity, where a tracklet could match multiple detections or vice versa. Ambiguity often arises from IoU-based matches when tracked objects are close, causing significant overlap in bounding boxes and similar IoU scores. If IoU is below the matching threshold for more than tracklet-detection pair of a tracklet or of a detection, we treat it as ambiguity.
@@
    We check if the considered tracklet’s mask is actually visible at the scene. Subjects can be entirely occluded resulting in no mask prediction at the current frame.
@@
    We check if the mask prediction is confident enough, i.e. if the average mask probability for the given object from the mask propagator is above the set mask confidence threshold.
Further, we compute two key ratios between the mask and detection bounding box:
@@
where $pix(\cdot)$ denotes pixels of the given instance and $mask(\cdot)$ denotes the mask assigned to the tracklet. ||$\cdot$|| ⋅ | denotes the cardinality of the set. Note that all $mm_{1},mm_{2}\in[0,1]$. In Fig. 4, we show how $mm_{1}$ and $mm_{2}$ can vary depending on mask and bounding box position. We discuss it more in detail in the supplementary material, Appendix F. Going further with the conditions:
@@
where $costs^{i,j}$ denotes the cost between tracklet i and detection j. With this fusion of the available information, we consider both modalities, masks and boxes to enhance the association process. The updated cost matrix, enriched by the mask signal, is passed to the Hungarian matching algorithm to find optimal tracklet-detection pairs.
Following conditions 1-4. ensures that the mask cue is controlled, and the cost matrix is updated only when the mask is reliable. The impact of each condition, along with the absence of any, is demonstrated in the ablation study (Sec. 4.3, Tab. 1). Further details on these conditions are included in the supplementary material, in Appendix F.
Further, using $mm_{1}$ directly to influence the cost matrix could be misleading, as multiple masks could fully fit within the same bounding box, all resulting in $mm_{1}=1.0$. Therefore, we use $mm_{1}$ only as a gating condition and $mm_{2}$ to influence the cost matrix.
Our baseline is optimized for bounding boxes, so we retain the use of the Hungarian matching algorithm over the cost matrix, but we carefully incorporate the mask signal to enhance the association process. In the ablation study (Sec. 4.3), we show that combining bounding boxes with the mask is more effective than using the mask alone, and demonstrate the advantages of incorporating the mask signal as an additional association cue over a purely bounding box-based approach.
In Figs. 1 and 2, we show that McByte can handle challenging scenarios such as long-term occlusion in crowd and association of ambiguous boxes due to the mask signal in the controlled manner as an association cue.
Figure 4: Cases showing the differences in $mm_{1}$ and $mm_{2}$ (Sec. 3.3) values of a mask (in blue) within a bounding box. The most optimal case for the mask to provide a good guidance is the second one from the left, where both $mm_{1}$ and $mm_{2}$ are as close to $1$ as possible.
@@ 3.4 Handling camera motion
When the camera moves, tracklet and detection bounding boxes may become less accurate due to object motion and blurring. To address this, we integrate camera motion compensation (CMC) into our process (which fuses mask and bounding box information) to enhance the accuracy of bounding box estimates. This approach follows existing methods [12, 27]. More details on CMC are provided in the supplementary material, in Appendix F. The final McByte version includes CMC, and its impact is shown in the ablation study in Sec. 4.3, Tab. 1.

### v3

# Temporally Propagated Masks and Bounding Boxes: Combining the Best of Both Worlds for Multi-Object Tracking

@@ Abstract
Multi-object tracking (MOT) involves identifying and consistently tracking objects across video sequences. Traditional tracking-by-detection methods, while effective, often require extensive tuning and lack generalizability. On the other hand, segmentation mask-based methods are more generic but struggle with tracking management, making them unsuitable for MOT. We propose a novel approach, McByte, which incorporates a temporally propagated segmentation mask as a strong association cue within a tracking-by-detection framework. By combining bounding box and propagated mask information, McByte enhances robustness and generalizability without per-sequence tuning. Evaluated on four benchmark datasets - DanceTrack, MOT17, SoccerNet-tracking 2022, and KITTI-tracking - McByte demonstrates performance gain in all cases examined. At the same time, it outperforms existing mask-based methods. Implementation code will be provided upon acceptance.
@@ 1 Introduction
In this paper, we explore using a temporally propagated segmentation mask as an association cue to assess its effectiveness in MOT. We propose a novel tracking-by-detection method that combines mask propagation and bounding boxes to improve the association between tracklets and detections. The mask propagation is managed according to the tracklet lifespan, while the mask is used in a controlled manner to enhance tracking performance. Since the temporal mask propagation model is trained on a large dataset, it makes the entire tracking process more generic. Unlike existing tracking-by-detection methods, our approach does not require tuning hyperparameters for each video sequence.
Tracking multiple objects at once often involves handling challenging occlusions, where only a small part of the subject might be visible, e.g. a leg of a person. Temporally propagated mask can be especially helpful in such cases, when the visible shape can considerably differ from the subject. A visual example is presented in Fig. 1.
We note explicitly that incorporating temporally propagated mask, which also involves temporal coherency, is different than using a static mask coming directly from an image segmentation model (such as SAM [20]) independently per each frame. We detail it in Sec. 3.2. To the best of our knowledge, using the mask temporal propagation within the problem of MOT has not been done before.
Our work covers specifically the problem of MOT, where we associate newly detected bounding boxes with previously tracked bounding boxes. Other challenges, such as Single Object Tracking (SOT), Video Object Segmentation (VOS), or Multi-Object Tracking and Segmentation (MOTS), although relevant, necessitate further modeling and are beyond the scope of this paper. In our work, we measure specifically the performance of bounding box association. For the mask performance, we refer the reader to the related works of temporal propagation models  [6, 8].
We evaluate our incorporation of the temporally propagated mask as an association cue against a baseline tracker, showing clear benefits for MOT. Our tracker is tested on four diverse MOT datasets, achieving the highest performance from tracking-by-detection algorithms on DanceTrack [33], MOT17 [28], SoccerNet-tracking 2022 [9] and KITTI-tracking [16]. These results highlight the advantages of using mask propagation, eliminating the need for per-sequence hyper-parameter tuning.
@@
    We propose a novel approach adapting a temporally propagated object segmentation mask as a strong and powerful cue, which, contrarily to a static mask, is incorporated into MOT for the first time. We describe the approach in Sec. 3.3.
@@
    We design an MOT tracking algorithm substantiating the idea and incorporating the temporally propagated mask as an association cue between tracklets and detections, detailed in Secs. 3.1 and 3.2. The tracker overcomes the limitations of mask-based approaches by performing proper tracklet management and including other important association cues as well as the limitations of the baseline tracking-by-detection approaches, by making the tracking process more robust and generic.
@@
    We evaluate and discuss the use of the temporally propagated mask as an association cue with different conditions within a tracking-by-detection algorithm in the ablation study (Sec. 4.3). Further, we evaluate our tracker incorporating the mask cue on four different datasets (in Sec. 4.4), comparing it to state-of-the-art tracking-by-detection approaches and demonstrating performance gain thanks to the attentive mask propagation usage, while not tuning per-sequence parameters.
Refer to caption Refer to caption Refer to caption
Figure 1: Temporally propagated mask can be helpful in cases of high occlusion. The person with the red mask is tracked only by its limited visible parts (pointed by yellow arrows for the clarity). Input image data from [28]. Best seen in color.
Figure 2: McByte tracking pipeline with the mask cue guidance. Temporally propagated mask signal is incorporated as an association cue in the tracklet-detection association steps.
@@ 2 Related Work
Different types of algorithms for tracking multiple objects at the same time have been developed in the community. We provide a brief overview of the existing methods related to our work.
2.1 Transformer-based and other types of method
Transformer-based methods [14, 41, 13, 38] use attention mechanisms to learn tracking trajectories and object associations from training data, following an end-to-end approach. These methods perform well on video sequences where subjects remain mostly on the scene, such as DanceTrack [33], but struggle with sequences where subjects frequently enter and exit, like in the MOT17 dataset [28], despite being designed for MOT. For this reason, we do not directly compare our method with transformer-based approaches in the main paper, though their results are listed together with the other approaches and ours in Appendix B.
Other types of MOT methods have also been proposed in the community, such as global optimization (offline) methods [5] or joint detection and tracking methods [39, 37]. Analogously to the transformer-based methods, they are not directly comparable and we list them for reference in Appendix B.
@@ 2.2 Tracking-by-detection methods
ByteTrack [40] is a powerful tracking-by-detection algorithm that uses the YOLOX [15] detector, associating tracklets based on the intersection over union (IoU) between bounding boxes of tracklets and new detections. It provides strong tracklet management and serves as a solid baseline for MOT. Several works have built on ByteTrack. OC-SORT [4] enhances state estimation by computing virtual trajectories during occlusion. StrongSORT [12] adds re-identification (re-ID) features as cues, camera motion compensation, and NSA Kalman filter [11]. C-BIoU [35] extends the association process by buffering (enlarging) bounding boxes, while HybridSORT [36] adds cues like confidence modeling and height-modulated IoU alongside existing strong cues.
Although these algorithms perform well on popular MOT datasets, they are highly sensitive to parameters. ByteTrack, for instance, explicitly tunes parameters like high-confidence detection thresholds per test sequence, as noted on its GitHub page [^1] https://github.com/ifzhang/ByteTrack, see: Test on MOT17. and in the code [^2] https://github.com/ifzhang/ByteTrack/blob/main/yolox/evaluators/mot_evaluator.py, lines 146-157 (at the moment of submitting this paper), which affects the tracking performance. Extensions of ByteTrack [12, 4, 27, 36], sharing the same code also rely on per-sequence parameter tuning. This tuning process is costly and impractical for larger datasets like DanceTrack [33], SoccerNet-tracking [9], and KITTI-tracking [16], as well as for generalizing across datasets. In contrast, our method, incorporating the temporally propagated segmentation mask as an association cue, avoids per-sequence tuning, making it more robust and generic, as we further demonstrate in Sec. 4.4.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Visual output comparison between the baseline and McByte. With the temporally propagated mask guidance, McByte can handle longer occlusion in the crowd - see the subject with ID 54 on the output of McByte. Input image data from  [28]. Best seen in color.
@@ 2.3 Segmentation mask models
Mask temporal propagation. XMem [6] is a mask temporal propagation model for video object segmentation (VOS), based on the Atkinson-Shiffrin Memory Model [1], enabling long-term tracking of segmentation masks. Its successor, Cutie [8], improves segmentation by incorporating object encoding from mask memory and better distinguishing the object from the background. Note that image segmentation models like SAM [20] only constitute the creation of the initial mask at the designated frames. Mask temporal propagation models then aim to infer the referred mask over the next frames in a video sequence.
While generally powerful in their target domains, both XMem and Cutie are not directly suitable for MOT as they do not involve bounding boxes and might provide inaccurate mask predictions as listed in their performance on video object segmentation [6, 8]. Therefore, we propose a novel MOT algorithm that combines temporally propagated masks with bounding boxes, improving tracking performance. Our ablation studies in Sec. 4.3 show that using the mask in a controlled manner, and along with other MOT mechanisms, is more effective than relying solely on the mask signal.
Mask-based tracking systems. Segmentation mask models have already been used to build tracking systems. DEVA [7], which enhances XMem [6], proposes decoupled video segmentation and bi-directional propagation, and builds a tracking system with boxes and masks. However, MOT requires a robust tracklet management and an occlusion handling system, which is not available within DEVA.
@@
Existing mask-based tracking algorithms, though trained on large datasets for extracting semantic features, fail to handle challenges like occlusion and track initiation. We discuss it more in detail in Sec. 4.5, where we show that their performance on MOT tasks is significantly lower than other methods. Our proposed MOT tracker incorporates the temporally propagated mask as an association cue while leveraging powerful mechanisms like tracklet management and additional cues (e.g., bounding box position, motion). As we demonstrate in Sec. 4.5, our approach performs considerably better than existing mask-based systems, even when they are adapted to MOT format and evaluation.
@@ 3.1 Preliminaries
In our work, we study a temporally propagated segmentation mask as a powerful association cue for MOT. We extend the baseline algorithm [40] and combine the mask information with bounding box information to create our novel masked-cued algorithm, which we call McByte. Fig. 2 shows the overview of our tracking pipeline pointing to where the temporally propagated mask is involved as an association cue. We note that we consider the problem of MOT, where we perform and evaluate the association between the bounding boxes of detections and tracklets.
In the next sections, we describe the creation and handling of the mask within our MOT tracking algorithm (Sec. 3.2) and our regulated use of the temporally propagated mask as an association cue (Sec. 3.3).
@@ 3.2 Mask creation and handling
Since the temporally propagated mask as a cue has to facilitate associating a tracklet with the correct detection, it must be well correlated with the tracklets and handled accordingly. This process is originally not straightforward and thus we design the following approach of mask handling to synchronise it with the processed tracklets.
During tracking in McByte, each tracklet gets its own mask, which is then temporally propagated across frames to update the mask predictions. At first, we use an image segmentation model to create a segmentation mask for each new tracklet. It is performed only for a newly appeared object and to initialize a new mask.
Separately, during the next frames, a temporal propagator is used to infer the updated mask positions, while aiming to keep the spatio-temporal consistency of the mask. The propagator predictions are analyzed and used in the tracklet-detection association process, as detailed in Sec. 3.3. We manage the temporally propagated masks in sync with the tracklet management system, creating new masks for new tracklets and removing them when a tracklet is terminated.
We provide more detailed information and a diagram including both mask creation and mask temporal propagation in Appendix E.
@@ 3.3 Regulated use of the mask
A temporally propagated (TP) segmentation mask can be a strong association cue if used correctly. However, mask predictions from the temporal propagator might sometimes be incorrect, as seen in related works [6, 8] and thus unreliable, making it essential to regulate the mask’s use as a cue.
In the association process, we update the cost matrix entries using the TP mask, particularly in cases of ambiguity, where a tracklet could match multiple detections or vice versa. Ambiguity often arises from IoU-based matches when tracked objects are close, causing significant overlap in bounding boxes and similar IoU scores. If IoU is below the matching threshold for more than one tracklet-detection pair, we treat it as ambiguity.
@@
    We check if the considered tracklet’s TP mask is actually visible on the scene. Subjects can be entirely occluded resulting in no mask prediction at the current frame.
@@
    We check if the TP mask prediction is confident enough, i.e. if the average mask probability for the given object from the mask propagator is above the set mask confidence threshold.
Further, we compute two key ratios between the TP mask and detection bounding box:
@@
where $pix(\cdot)$ denotes pixels of the mask or within the bounding box, and $mask(\cdot)$ denotes the TP mask assigned to the tracklet. ||$\cdot$|| ⋅ | denotes the cardinality of the set. Note that all $mm_{1},mm_{2}\in[0,1]$. In Fig. 4, we show how $mm_{1}$ and $mm_{2}$ can vary depending on the TP mask and bounding box position. We discuss it more in detail in Appendix E. Going further with the conditions:
@@
where $costs^{i,j}$ denotes the cost between tracklet i and detection j. With this fusion of the available information, we consider both modalities, TP masks and bounding boxes to enhance the association process. The updated cost matrix, enriched by the mask signal, is passed to the Hungarian matching algorithm to find optimal tracklet-detection pairs.
Following conditions 1-4. ensures that the TP mask cue is controlled, and the cost matrix is updated only when the mask is reliable. The impact of each condition, along with the absence of any, is demonstrated in the ablation study (Sec. 4.3, Tab. 1). Further details on these conditions are included in Appendix E.
Further, using $mm_{1}$ directly to influence the cost matrix could be misleading, as multiple TP masks could fully fit within the same bounding box, all resulting in $mm_{1}=1.0$. Therefore, we use $mm_{1}$ only as a gating condition and $mm_{2}$ to influence the cost matrix.
Our baseline is optimized for bounding boxes, so we retain the use of the Hungarian matching algorithm over the cost matrix, but we carefully incorporate the TP mask signal to enhance the association process. In the ablation study (Sec. 4.3), we show that combining bounding boxes with the mask is more effective than using the mask alone, and demonstrate the advantages of incorporating the TP mask signal as an additional association cue over a purely bounding box-based approach.
In Fig. 3, we show that McByte can handle challenging scenarios such as long-term occlusion in crowd due to the TP mask signal in the controlled manner as an association cue. More visual examples are available in Appendix E.
Figure 4: Cases showing the differences in $mm_{1}$ and $mm_{2}$ (Sec. 3.3) values of a temporally propagated mask (in blue) within a bounding box. The most optimal case for the mask to provide a good guidance is the second one from the left, where both $mm_{1}$ and $mm_{2}$ are as close to $1$ as possible.
@@ 3.4 Handling camera motion
When the camera moves, tracklet and detection bounding boxes may become less accurate due to object motion and blurring. To address this, we integrate camera motion compensation (CMC) into our process (which fuses temporally propagated mask and bounding box information) to enhance the accuracy of bounding box estimates. This approach follows existing methods [12, 27]. More details on CMC are provided in Appendix E. The final McByte version includes CMC, and its impact is shown in the ablation study in Sec. 4.3, Tab. 1.